OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!! Input the equation of the given line in standard form. The line with m = and the point (-2, -3).

OpenStudy (anonymous):

the standard form is y=mx+b

OpenStudy (anonymous):

for thisinformation u can use this equation \[y - y _{1}=m \times (x - x _{1})\]

OpenStudy (anonymous):

@blahblahblah375 ok?

OpenStudy (anonymous):

yeah but i keep getting y=1/2x-7 but thats not the answer

OpenStudy (anonymous):

@sorour

OpenStudy (anonymous):

m=?

OpenStudy (anonymous):

@blahblahblah375 ,m=?

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

y=-1/2x-4

OpenStudy (anonymous):

y+3=-1/2(x+2) y=-1/2x-1-3 y=-1/2x-4

OpenStudy (anonymous):

@blahblahblah375 , ok?

OpenStudy (anonymous):

the equation they set up is